Tonycoolkid21:

what can divide 100 and 125

AZ:

Can you list the factors of 100 and 125? There are a lot of numbers that can divide both 100 and 125

Tonycoolkid21:

15 20 10 25

AZ:

Let's start with smaller numbers and work our way up What are the factors of 100 basically what numbers multiply together to give you 100 2 * 50 = 100 but what else?

AZ:

@tonycoolkid21 wrote:
15 20 10 25
are those answer choices or were you trying to answer my question? o.O

Tonycoolkid21:

yea

AZ:

okay, so take a calculator and 100 / 15 = ?? is a whole number or what about 100/20 = ?? and 100/10 = ?? 100 / 25 = ??

AZ:

and then similarly 125 / 15 = ?? 125 / 20 = ?? 125 / 10 = ?? 125 / 25 = ??

MASTERKIY:

Whelllll I think of it as, 100 as 1.00, and then how many quarters make a dollar, then what is the value of them coins. That would make my answer~

Tonycoolkid21:

ok so 100/15 = 6

AZ:

no it doesn't

Tonycoolkid21:

0.15

Tonycoolkid21:

i did not do anything

AZ:

which one gives you just a whole number like 100/10 = 10 but 125/ 10 = 12.5 12.5 is not a whole number

AZ:

so like similarly divide for the other numbers 100 / 20 = ?? 125 / 20 = ?? 100/ 25 = ?? 125/ 25 = ??

Tonycoolkid21:

25

AZ:

there you go!

AZ:

Good job :)

MASTERKIY:

Bro- all that work you should have thought of it as moneyyy best way n it makes u feel broke

Tonycoolkid21:

wdym

MASTERKIY:

If you thought of it as money u would get the same answer

AZ:

Maybe for this particular question? But it wouldn't always work like that

MASTERKIY:

Thats what im referring to, this question could have been solved fast thinking of it as money. Could of saved a few extra minutes for a stretch or sum

AZ:

What if the two numbers you get are 111 and 33 how do you use money to help you find the GCF

MASTERKIY:

brevo- again I was referring to THIS particular question. :) not for you to be confused.